When a plane mirror on which a ray is incident is rotated through 100, the reflected ray is rotated through what angle
1
Expert's answer
2020-08-12T16:18:56-0400


Let the initial angle of incident be "\angle AOB = \alpha". Then the initial reflection angle also will be: "\angle EOA = \alpha". Suppose we rotated the mirror by the angle "\angle AOC = \theta = 100 \degree". Then, according to the figure, we need to fing the angle "\angle EOD" - the angle between old and new reflected rays. It will be:


"\angle EOD = \angle EOA + \angle AOB + \angle BOC + \angle COD"

First two angles we've already known, so let's fihd the last two.


"\angle BOC = \angle AOC - \angle AOB = \theta-\alpha = 100\degree-\alpha"

As incident and reflection angles:


"\angle COD = \angle BOC = 100\degree - \alpha"

Thus, obtain:


"\angle EOD = \alpha + \alpha + (100\degree - \alpha) + (100\degree - \alpha) = 200\degree"

Answer. "200\degree".